OpenStudy (anonymous):

Am I doing this right? 4x + y + 11 5x + 3y = 1 change my equation y = 4x + 11 5x + 3( 4x + 11) =1 5x 12 + 33 =1 combine like terms -7x 33 =1 33 33 -7x 32 7x 7 4 plug for in the 3 equation y = 4 (4) + 11 16 16 5?

OpenStudy (anonymous):

Um, your first equation is incorrect..

OpenStudy (anonymous):

You put two + signs..

OpenStudy (anonymous):

Is it this? 4x + y = 11

OpenStudy (anonymous):

u have to take 4x from both sides so one would be -4x which will make it y = 11 4x?

OpenStudy (anonymous):

my bad that is y = -4x +11

OpenStudy (anonymous):

Yes, that's correct.

OpenStudy (anonymous):

y = -4x + 11 is correct. But you did something wrong. -7x + 33 = 1 Subtract 33 to both sides, 1 - 33 = ?